OpenStudy (anonymous):

The sum of the measures of the interior angles of a convex polygon is 3060º. How many sides does the polygon have? A. 15 B. 17 C. 18 D. 19

OpenStudy (agl202):

180(n-2) = 3060 n-2 = 3060/180 n-2 = 17 n = 19 sides Hope this helps u!!! @jakematt24

OpenStudy (anonymous):

Thank You @Agl202

OpenStudy (agl202):

=)